OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 69 = x + 1/2= (x - 6) What is the regular price of a child's ticket? (4 points) $38 $42 $48 $66

OpenStudy (chris911):

$42 refer to https://answers.yahoo.com/question/index?qid=20110707171138AAHOgcp

OpenStudy (anonymous):

Awesome thanks

OpenStudy (chris911):

as always ur welcome ;)

OpenStudy (nurali):

69 = x + 1/2(x - 6) 69 = x + (x / 2) - (6 / 2) 69 = 3x/2 - 3 69 + 3 = 3x / 2 72 = 3x / 2 72 X 2 = 3x 144 = 3x 144 / 3 = x 48 = x x = 48

OpenStudy (chris911):

darn

OpenStudy (anonymous):

It was 42

OpenStudy (chris911):

XD haha Nurali I did it again XD lol jk jk